A flywheel in the form of a uniformly thick disk of radius 1.73 m has a mass of 28.6 kg and spins counterclockwise at 283 rpm . Calculate the constant torque required to stop it in 2.25 min .
Answer #1

radius of the disk r = 1.73 m

mass m = 28.6 kg

No. of revolution N = 283 rpm

angular speed w = 2N

w = 2 x 283 rad/min

= 2 x 283/60 rad/s

w = 29.64 rad/s

since flywheel stops in 2.25 min (2.25 x 60 = 135 s) , using equation of rotational motion

wf = wi + t

0 = 29.64 + x 135

angular acceleration = -0.22 rad/s2

also we know from Newtons 2nd law

torque = I x

= (1/2) x m x r2 x

= 0.5 x 28.6 x (1.73)2 x 0.22

= 9.39 N-m